Break-Even Units, Contribution Margin Ratio, Margin of Safety

Khumbu Company's projected profit for the coming year is as follows:

Total Per Unit
Sales $1,275,000 $25.50
Total variable cost 459,000 9.18
Contribution margin $ 816,000 $ 16.32
Total fixed cost 518,670
Operating income $ 297,330

Required:

1. Compute the break-even point in units. If required, round your answer to nearest whole value.
units

2. How many units must be sold to earn a profit of $240,000? If required, round your answer to nearest whole value.
units

3. Compute the contribution margin ratio. If required, round your answer to nearest whole number.
%

Using the rounded ratio from above, compute the additional profit that Khumbu would earn if sales were $160,000 more than expected.
$

4. For the projected level of sales, compute the margin of safety in units.
units
Answer #1

1) Break even unit = Fixed cost/Contribution margin per unit = 518670/16.32 = 31781 units

2) Required unit = (518670+240000)/16.32 = 46487 Units

3) Contribution margin ratio = 16.32/25.50 = 64%

Additional profit = 160000*64% = 102400

4) Margin of safety in units = 50000-31781 = 18219 Units